Determine whether ~u and ~v are orthogonal vectors, make an acute or obtuse angle: (1.1) ~u =< 1, 3, −2 >, ~v =< −5, 3, 2 >. (2) (1.2) ~u =< 1, −2, 4 >, ~v =< 5, 3, 7 >.


1
Expert's answer
2021-07-01T12:27:31-0400

Let us use the dot products of vectors.


(1.1) Since "\overline u \cdot \overline v =< 1, 3, −2 >\cdot< −5, 3, 2 >=-5+9-4=0," we conclude that the vectors are orthogonal.


(1.2) Since "\overline u \cdot \overline v=< 1, −2, 4 >\cdot< 5, 3, 7 >=5-6+28=27>0," we conclude that the vectors

make an acute angle.